OpenStudy (anonymous):

By using the fact that lim x->0 (sinx/x) = 1,or otherwise find lim x->0 (2x^3+4x)/(sin2x) PLEASE HELP

OpenStudy (anonymous):

So you don't get to use l'Hospital's rule?

OpenStudy (anonymous):

No I don't.

OpenStudy (anonymous):

Could you please assist me? The help would be very much appreciated.

OpenStudy (anonymous):

\[ \sin(2x) = 2\sin(x)\cos(x) \]

OpenStudy (anonymous):

\[ 2x^3+4x =2x(x^3+2) \]

OpenStudy (anonymous):

then just factor out the \(\sin(x)\) and \(x\)

OpenStudy (anonymous):

Since \[ \lim f(x)g(x) = [\lim f(x)][\lim g(x)] \]

OpenStudy (anonymous):

do you get it?

OpenStudy (anonymous):

it would be x/sinx * (x^3+2)/cosx?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

so the answer is 2?

OpenStudy (anonymous):

Oh thank you, wow how do I like "pay" you on this site?

OpenStudy (anonymous):

Become a fan?